# {tocentry}[Uncaptioned image]

### v2

@@ 0.1 Introduction
In this work, we derive a generalized coupled cluster theory (GCC) that avoids all the unphysical behaviors mentioned above. The complications related to matrix defects in the non-Hermitian eigenvalue problem can be handled using the techniques developed in similarity-constrained coupled cluster theory and will not be discussed here. The application of GCC with singles and doubles excitations (GCCSD) will be illustrated for several molecular systems where the coupled cluster singles and doubles (CCSD) model fails to give a correct description. We will also formulate the GCC2 model, which is a generalization of the well-established CC2 model.19 Finally, we will propose a procedure to eliminate bifurcations in Hartree-Fock and density functional theory, which in their present formulation are unable to describe ground state conical intersections.20, 21
@@ 1 Generalized coupled cluster theory
where eq. 4 is the amplitude equations and $E_{0}$ is the coupled cluster energy.22 The similarity-transformed Hamiltonian is given by $\bar{H}=\exp{(-T)}H\exp{(T)}$, where $H$ is the electronic Born-Oppenheimer Hamiltonian.
@@
where $\Omega_{k}=(\mathbf{S}^{-1}\Omega)_{k}$, $\Delta t_{k}=(\mathbf{S}^{-1}\Delta t)_{k}$, $\tau_{k}=(\tau^{T}\mathbf{S})_{k}$, and $\omega_{k}$ is an eigenvalue of $\mathbf{A}$. It is clear that if $\omega_{k}$ is close to zero, then the higher-order terms in the expansion dominate, giving rise to a bifurcation into more than one solution. Depending on the properties of these higher-order terms, we may obtain different situations with two or more real solutions or even none. Before we discuss our solution to the bifurcation problem, let us consider the equation of motion eigenvalue problem23, 24 for $\bar{H}$
@@ 2 Scaling with system size
Size extensivity of total energies and size intensivity of excitation energies are two essential properties of coupled cluster theory that ensure scalability to large systems without losing accuracy.25, 22 Therefore, we will show that these properties are maintained in GCC. We consider a system composed of two non-interacting subsystems A and B. The total Hamiltonian is the sum of two separate terms $H=H_{A}+H_{B}$. When the cluster amplitudes are extensive, $T=T_{A}+T_{B}$, we have that $\bar{H}=\bar{H}_{A}+\bar{H}_{B}$, where $\bar{H}_{X}=\exp(-T_{X})H_{X}\exp(T_{X})$. The Hartree-Fock reference is given by the direct product state $|\mathrm{HF}\rangle=|\mathrm{HF_{A}}\rangle\otimes|\mathrm{HF_{B}}\rangle$ and the excitation manifold is ordered as $\{\langle\mu_{A}|,\langle\mu_{B}|,\langle\mu_{AB}|\}$.
It is well known that when the cluster amplitudes are extensive then the right eigenvectors of the Jacobian in eq. 21 are intensive. 25 Thus we consider a right eigenvector located in system A and denote the excitation operator $R^{A}_{1}$. The corresponding left operator $L^{A}_{1}$ does not have components in B but can have non-zero elements in the AB part of the operator. We first note that the amplitude equations in eq. 20 take the form
@@ 3 Applications
The GCCSD framework, for an arbitrary number of projected states, was implemented in a local development version of the $e^{T}$ program 26 and used in all the calculations reported here. Initially, we will consider the lithium fluoride molecule, a standard benchmark system for multireference coupled cluster methods.27 At a bond distance of about $6.75$ the molecule has an avoided crossing 28 between the $2^{1}\Sigma^{+}$ state and the ground state $1^{1}\Sigma^{+}$. We employ an aug-cc-pVDZ basis and find two solutions to the CCSD ground state equations. The CCSD solution (Fig. 1b), obtained when starting at distances close to the equilibrium bond length ($1.5$), shows an unexpected increasing diverging trend in the energies of all 3 states when extended past the avoided crossing and restarting the algorithm from the previous point. Another CCSD solution is found (Fig. 1c) when starting at a long bond distance ($9.0$). We observe a decreasing diverging trend in the energies when extended to shorter distances. The solution eventually changes discontinuously to the other CCSD solution at about $2.0$. On the other hand, the GCCSD curves shown in Fig. 1a are continuous for the whole range of bond distances. Thus, GCCSD bridges the CCSD solution for distances shorter than the avoided crossing with the other CCSD solution beyond the avoided crossing point. In doing so, GCCSD corrects the divergent behavior seen in the two individual CCSD solutions.
Figure 2: The GCCSD potential energy surfaces of $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ in ethylene. The basis is aug-cc-pVDZ and the energies are plotted relative to the average energy $\frac{1}{2}(E_{0}+E_{1})$. The $\varepsilon$-MECI structure is shown in the middle, which corresponds to the geometry at the point $(\mathrm{g,h})=(0.0,0.0)$.
Figure 3: The CCSD potential energy surfaces of $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ in ethylene. The basis is aug-cc-pVDZ and the energies are plotted relative to the average energy $\frac{1}{2}(E_{0}+E_{1})$. There are three notable regions: in A, a mismatch in energies appears due to the phase effect; in B, a new set of flipped solutions is obtained, characterized by negative excitation energies; and in C, the region where we were not able to converge the ground state equations.
We now consider a ground state conical intersection in ethylene. We will start from the CCSD $\varepsilon$-MECI using an aug-cc-pVDZ basis to explore the potential energy surfaces close to a conical intersection. The $\varepsilon$-MECI structure is reported by Angelico et al. 18 and is shown in Fig. 2. We further employ the $\mathbf{g}$ and $\mathbf{h}$ vectors calculated at the $\varepsilon$-MECI geometry using the CCSD algorithm described in Ref.29. Further computational details are reported in the Supporting Information.
@@
In concluding our investigation of ethylene we will discuss the convergence properties of the framework. We use the ethylene A geometry from the extensivity study above, given in the Supporting Information. In Table 2 we report the energies for the ground state and the two first excited states. The reduced space numbers are almost identical to the full space and in turn, these are similar to the CCSD numbers. Even though the number of iterations indicates a factor of $2.5$ in computational cost between GCCSD and CCSD, the wall times show only a factor of $1.7$. This is due to relatively fewer linear transformations in GCCSD compared to CCSD per iteration.
To demonstrate the GCCSD method when projecting out several states we consider the $\mathrm{S_{1}}/\mathrm{S_{2}}$ conical intersection in thymine. We project out the eigenvectors for the two lowest eigenvalues of the Jacobian. This means the reduced space matrix is a $3\times 3$ matrix. We use an initial structure together with $\mathbf{g}$ and $\mathbf{h}$ vectors that were determined in Ref. 12. These are reported in the Supporting Information. In Fig. 6 we show the $\mathrm{S}_{1}$ and $\mathrm{S}_{2}$ potential energy surfaces for CCSD and GCCSD. Both surfaces have a defect, that can be removed using a similarity-constrained transformation, and they look very similar. The center of the defect is slightly shifted when comparing the two.
Figure 6: CCSD (a) and GCCSD (b) potential energy surfaces of $\mathrm{S}_{1}$ and $\mathrm{S}_{2}$ in thymine with cc-pVDZ basis. All energies are plotted in eV, relative to the average $\frac{1}{2}(E_{1}+E_{2})$. The conical intersection structure is shown in the middle, which corresponds to the geometry at the point $(\mathrm{g,h})=(0.0,0.0)$.
As a final example, we show a $\mathrm{S}_{0}$/$\mathrm{S}_{1}$ intersection in 2,4-cyclohexadien-1-ylamine. The geometry was obtained from the database in Ref. 30 and the structure is shown in Fig. 7. The $\mathbf{g}$ and $\mathbf{h}$ vectors are calculated using the CCSD algorithm described in Ref.29. In the same figure, we show the two potential energy surfaces in two different representations. The conical shape of the intersection is visible in both plots.
Figure 7: The GCCSD potential energy surfaces of $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ in 2,4-cyclohexadien-1-ylamine with cc-pVDZ. In (a) the energies are plotted in eV relative to the average energy $\frac{1}{2}(E_{0}+E_{1})$ whereas in (b) the total energies are shown in Hartree. The initial structure is shown in the middle, which corresponds to the geometry at the point $(\mathrm{g,h})=(0.0,0.0)$.
@@ 4 Further perspectives
The framework presented above can be extended to other electronic structure methods with similar problems accounting for the geometric phase associated with conical intersections with the ground state. For instance, time-dependent Hartree-Fock (TDHF) and time-dependent density functional theory (TDDFT) are unable to describe the intersection as the ground and excited states are decoupled.20, 31 Other methods such as the algebraic diagrammatic construction (ADC) hierarchy32 also do not include coupling to the ground state.
@@
The entire coupled cluster hierarchy is in general affected by the geometric phase issues highlighted by Kjønstad and Koch.13 Here we explicitly mention the CC2 method,19 which is frequently used for calculating excitation energies of large molecules, due to the favorable balance between computational cost and accuracy compared to CCSD. The CC2 model is viewed as the best alternative to second-order Møller-Plesset theory as both ground and excited states are available with the same computational cost, and would therefore be an excellent candidate for nonadiabatic dynamics. However, the CC2 method also fails to describe ground state conical intersections, as shown in Figure 8 for the ethylene molecule. Using the GCC framework above we may extend CC2 to GCC2 by considering the CC2 Hamiltonian matrix
@@
The explicit expressions for $\mathbf{\Omega}^{\mathrm{CC2}}$ and $\mathbf{A}^{\mathrm{CC2}}$ can be found in Ref.19. We will report the implementation of GCC2 elsewhere together with a detailed benchmarking of the method.
We now outline how the conical intersection problem could be solved for closed-shell Hartree-Fock. Due to the closed shell form of the Hartree-Fock wave function a rotation of the orbitals cannot change the overall sign of the wave function. This implies that Hartree-Fock theory cannot detect an intersection with the excited state. However, as discussed in detail by Helgaker et. al,22 the eigenvalues of the electronic Hessian may become small and this can lead to bifurcations in the Hartree-Fock solution. Using the ideas from GCC we may remove these bifurcations. For this purpose, we parameterize the wave function in terms of orbital rotations
@@
where $P_{ai,bj}$ permutes $ai$ and $bj$ indices. We may now remove the Hessian eigenvector $\mathbf{r}$ from the $\bm{\kappa}$ vector, which is associated with a small eigenvalue that gives rise to bifurcations. This leads to a coupled set of equations that must be solved, in the same way as for GCC. The removed component may be included through diagonalization of the matrix
@@
In passing, we note that as the eigenvectors of the Hessian in eq. 40 are size-intensive then the resulting wave functions will be extensive. They can therefore serve as a reference for a coupled cluster model that extends the GCC framework presented here. Needless to say, the framework can also be used for Tamm-Dancoff TDDFT.21, 20

### v3

@@ 0.1 Introduction
Several approaches have been developed to circumvent the ground state intersection problem in electronic structure methods. These methods mainly involve starting from a different state of the system, for instance, a high-spin triplet state followed by a spin-flip (SF) in the equation-of-motion (EOM) treatment (SF-TD-DFT 19 and EOM-SF-CCSD 20) or starting from a double ionized state and adding electrons back in via the EOM framework (hh-TDA 21, 22 and DEA-EOM-CCSD 23, 24). Other methods, notably in density functional theory (DFT), have been developed that either change the kernel to account for double excited states 25 or by adding an additional double excited configuration to the Hermitian Tamm-Dancoff eigenvalue problem (TDDFT-1D 26). Ensemble DFT (eDFT) methods have also been applied to strongly correlated systems and extended to treat excited states (SI-SA-REKS 27). More recently, Schmerwitz et al. 28 developed an algorithm to determine saddle points in DFT methods that can be identified with electronic excited states, giving access to ground state intersections. This method is related to the norm-extended optimization scheme for multiconfigurational wave functions 29 that also determine saddle point states that are nonorthogonal.
In this work, we derive a generalized coupled cluster theory (GCC) that avoids all the unphysical behaviors mentioned above. The framework corrects the ground state wave function parametrization and does not consider a different state of the system as a starting point. The complications related to matrix defects in the non-Hermitian eigenvalue problem can be handled using the techniques developed in similarity-constrained coupled cluster theory and will not be discussed here. The application of GCC with singles and doubles excitations (GCCSD) will be illustrated for several molecular systems where the coupled cluster singles and doubles (CCSD) model fails to give a correct description. We will also formulate the GCC2 model, which is a generalization of the well-established CC2 model.30 Finally, we will propose a procedure to eliminate bifurcations in Hartree-Fock and density functional theory, which in their present formulation are unable to describe ground state conical intersections.31, 32
@@ 1 Generalized coupled cluster theory
where eq. 4 is the amplitude equations and $E_{0}$ is the coupled cluster energy.33 The similarity-transformed Hamiltonian is given by $\bar{H}=\exp{(-T)}H\exp{(T)}$, where $H$ is the electronic Born-Oppenheimer Hamiltonian.
@@
where $\Omega_{k}=(\mathbf{S}^{-1}\Omega)_{k}$, $\Delta t_{k}=(\mathbf{S}^{-1}\Delta t)_{k}$, $\tau_{k}=(\tau^{T}\mathbf{S})_{k}$, and $\omega_{k}$ is an eigenvalue of $\mathbf{A}$. It is clear that if $\omega_{k}$ is close to zero, then the higher-order terms in the expansion dominate, giving rise to a bifurcation into more than one solution. Depending on the properties of these higher-order terms, we may obtain different situations with two or more real solutions or even none. Before we discuss our solution to the bifurcation problem, let us consider the equation of motion eigenvalue problem34, 35 for $\bar{H}$
@@
We note that solving the eigenvalue problem for both the reduced matrix and the full matrix can sometimes result in a complex pair of eigenvalues and eigenvectors, which is what is observed in a small region close to the conical intersection (see Fig. 2 and Fig. S4). This situation is discussed in Ref. 7 for the case of conical intersections among excited states, but the same conclusions can be applied to the problem in Eq. 22. This final step presents the same characteristics as EOM-CC, where the partial ground state solution obtained in the reduced parameter space acts as the ground state, and the Jacobian matrix is replaced by the full space matrix.
@@ 2 Scaling with system size
Size extensivity of total energies and size intensivity of excitation energies are two essential properties of coupled cluster theory that ensure scalability to large systems without losing accuracy.36, 33 Therefore, we will show that these properties are maintained in GCC. We consider a system composed of two non-interacting subsystems A and B. The total Hamiltonian is the sum of two separate terms $H=H_{A}+H_{B}$. When the cluster amplitudes are extensive, $T=T_{A}+T_{B}$, we have that $\bar{H}=\bar{H}_{A}+\bar{H}_{B}$, where $\bar{H}_{X}=\exp(-T_{X})H_{X}\exp(T_{X})$. The Hartree-Fock reference is given by the direct product state $|\mathrm{HF}\rangle=|\mathrm{HF_{A}}\rangle\otimes|\mathrm{HF_{B}}\rangle$ and the excitation manifold is ordered as $\{\langle\mu_{A}|,\langle\mu_{B}|,\langle\mu_{AB}|\}$.
It is well known that when the cluster amplitudes are extensive then the right eigenvectors of the Jacobian in eq. 21 are intensive. 36 Thus we consider a right eigenvector located in system A and denote the excitation operator $R^{A}_{1}$. The corresponding left operator $L^{A}_{1}$ does not have components in B but can have non-zero elements in the AB part of the operator. We first note that the amplitude equations in eq. 20 take the form
@@ 3 Applications
The GCCSD framework, for an arbitrary number of projected states, was implemented in a local development version of the $e^{T}$ program 37 and used in all the calculations reported here. Initially, we will consider the lithium fluoride molecule, a standard benchmark system for multireference coupled cluster methods.38 At a bond distance of about $6.75$ the molecule has an avoided crossing 39 between the $2^{1}\Sigma^{+}$ state and the ground state $1^{1}\Sigma^{+}$. We employ an aug-cc-pVDZ basis and find two solutions to the CCSD ground state equations. The CCSD solution (Fig. 1b), obtained when starting at distances close to the equilibrium bond length ($1.5$), shows an unexpected increasing diverging trend in the energies of all 3 states when extended past the avoided crossing and restarting the algorithm from the previous point. Another CCSD solution is found (Fig. 1c) when starting at a long bond distance ($9.0$). We observe a decreasing diverging trend in the energies when extended to shorter distances. The solution eventually changes discontinuously to the other CCSD solution at about $2.0$. On the other hand, the GCCSD curves shown in Fig. 1a are continuous for the whole range of bond distances. Thus, GCCSD bridges the CCSD solution for distances shorter than the avoided crossing with the other CCSD solution beyond the avoided crossing point. In doing so, GCCSD corrects the divergent behavior seen in the two individual CCSD solutions.
Figure 2: (a) The GCCSD potential energy surfaces of $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ in ethylene. (b) A detailed view of the region close to the conical intersection. The basis is aug-cc-pVDZ and for each point the energies are plotted relative to the average energy $\frac{1}{2}(E_{0}+E_{1})$. A plot of the same region, showing the absolute energies of the two states in Hartree can be found in the Supporting Information. The $\varepsilon$-MECI structure is shown in the middle, which corresponds to the geometry at the point $(\mathrm{g,h})=(0.0,0.0)$.
Figure 3: The CCSD potential energy surfaces of $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ in ethylene. The basis is aug-cc-pVDZ and for each point the energies are plotted relative to the average energy $\frac{1}{2}(E_{0}+E_{1})$. There are three notable regions: in A, a mismatch in energies appears due to the phase effect; in B, a new set of flipped solutions is obtained, characterized by negative excitation energies; and in C, the region where we were not able to converge the ground state equations.
We now consider a ground state conical intersection in ethylene. We will start from the CCSD $\varepsilon$-MECI using an aug-cc-pVDZ basis to explore the potential energy surfaces close to a conical intersection. The $\varepsilon$-MECI structure is reported by Angelico et al. 18 and is shown in Fig. 2. We further employ the $\mathbf{g}$ and $\mathbf{h}$ vectors calculated at the $\varepsilon$-MECI geometry using the CCSD algorithm described in Ref.40. Further computational details are reported in the Supporting Information.
@@
In concluding our investigation of ethylene, we discuss the convergence properties of the framework. We use the ethylene A geometry from the extensivity study above, given in the Supporting Information. In Table 2 we report the energies for the ground state and the two first excited states. The reduced space numbers are almost identical to the full space and in turn, these are similar to the CCSD numbers. Even though the number of iterations indicates a factor of $2.5$ in computational cost between GCCSD and CCSD, the wall times show only a factor of $1.7$. This is due to relatively fewer linear transformations in GCCSD compared to CCSD per iteration.
To demonstrate the GCCSD method when projecting out several states we consider the $\mathrm{S_{1}}/\mathrm{S_{2}}$ conical intersection in thymine. We project out the eigenvectors for the two lowest eigenvalues of the Jacobian. This means the reduced space matrix is a $3\times 3$ matrix. We use an initial structure together with $\mathbf{g}$ and $\mathbf{h}$ vectors that were determined in Ref. 12. These are reported in the Supporting Information. In Fig. 6 we show the $\mathrm{S}_{1}$ and $\mathrm{S}_{2}$ potential energy surfaces for CCSD and GCCSD. The surfaces look very similar and they both have a defect that can be removed using a similarity-constrained transformation, as shown in Fig. S3 for CCSD. The center of the defect is slightly shifted when comparing the two.
Figure 6: CCSD (a) and GCCSD (b) potential energy surfaces of $\mathrm{S}_{1}$ and $\mathrm{S}_{2}$ in thymine with cc-pVDZ basis. All energies are plotted in eV, relative to the average $\frac{1}{2}(E_{1}+E_{2})$ for each point. The conical intersection structure is shown in the middle, which corresponds to the geometry at the point $(\mathrm{g,h})=(0.0,0.0)$.
As a final example, we show a $\mathrm{S}_{0}$/$\mathrm{S}_{1}$ intersection in 2,4-cyclohexadien-1-ylamine. The geometry was obtained from the database in Ref. 41 and the structure is shown in Fig. 7. The $\mathbf{g}$ and $\mathbf{h}$ vectors are calculated using the CCSD algorithm described in Ref.40. In the same figure, we show the two potential energy surfaces in two different representations. The conical shape of the intersection is visible in both plots. Again, when zooming in on the intersection region, a small defective area is observed (see Fig. S2 in the Supporting Information).
Figure 7: The GCCSD potential energy surfaces of $\mathrm{S}_{0}$ and $\mathrm{S}_{1}$ in 2,4-cyclohexadien-1-ylamine with cc-pVDZ. In (a) the energies for each point are plotted in eV relative to the average energy $\frac{1}{2}(E_{0}+E_{1})$ whereas in (b) the total energies are shown in Hartree. The initial structure is shown in the middle, which corresponds to the geometry at the point $(\mathrm{g,h})=(0.0,0.0)$.
@@ 4 Further perspectives
The framework presented above can be extended to other electronic structure methods with similar problems accounting for the geometric phase associated with conical intersections with the ground state. For example, time-dependent Hartree-Fock (TDHF) and time-dependent density functional theory (TDDFT) cannot describe the intersection because the ground and excited states are decoupled.31, 42 Other methods such as the algebraic diagrammatic construction hierarchy (ADC)43, 44 also do not include coupling to the ground state. The ADC framework relies on Møller-Plesset perturbation theory for the ground state wave function and as shown in Ref. 13 the perturbation series converge to an excited state close to the conical intersection. Recently, Taylor et al. 45 discussed the failure of the ADC(2) model at ground state intersections.
@@
The entire coupled cluster hierarchy is generally affected by the geometric phase issues highlighted in Ref. 13. Here we explicitly mention the CC2 method,30 which is frequently used for calculating excitation energies of large molecules, due to the favorable balance between computational cost and accuracy compared to CCSD. The CC2 model is viewed as the best alternative to second-order Møller-Plesset theory as both ground and excited states are available with the same computational cost and would therefore be an excellent candidate for nonadiabatic dynamics. However, the CC2 method also fails to describe ground state conical intersections, as shown in Figure 8 for the ethylene molecule. Using the GCC framework above, we may extend CC2 to GCC2 by considering the CC2 Hamiltonian matrix
@@
The explicit expressions for $\mathbf{\Omega}^{\mathrm{CC2}}$ and $\mathbf{A}^{\mathrm{CC2}}$ can be found in Ref.30. We will report on the implementation of GCC2 elsewhere together with a detailed benchmarking of the method.
We now outline how the conical intersection problem could be solved for closed-shell Hartree-Fock. Due to the closed shell form of the Hartree-Fock wave function, a rotation of the orbitals cannot change the overall sign of the wave function. This implies that Hartree-Fock theory cannot detect an intersection with the excited state. However, as discussed in detail by Helgaker et. al,33 the eigenvalues of the electronic Hessian may become small and this can lead to bifurcations in the Hartree-Fock solution. Using the ideas from GCC we may remove these bifurcations. For this purpose, we parameterize the wave function in terms of orbital rotations
@@
where $P_{ai,bj}$ permutes $ai$ and $bj$ indices. We may now remove the Hessian eigenvector $\mathbf{r}$ from the $\bm{\kappa}$ vector, which is associated with a small eigenvalue that gives rise to bifurcations. This leads to a coupled set of equations that must be solved in the same way as for GCC. The removed component may be included through diagonalization of the matrix
@@
In passing, we note that as the eigenvectors of the Hessian in eq. 40 are size-intensive then the resulting wave functions will be extensive. They can therefore serve as a reference for a coupled cluster model that extends the GCC framework presented here. Needless to say, the framework can also be used for Tamm-Dancoff TDDFT.32, 31